oosuweiuhfq:

954/0.9 please explain

surjithayer:

\[\frac{ 954 }{0.9 }=\frac{ 954 }{ \frac{ 9 }{ 10 } }=954\times \frac{ 10 }{ 9}=106\times10=1060\]